Find a number such that 6 more than one-half the number is two-thirds the number.
Answer by checkley77(12844)

x/2+6=2/3x
x/2-2x/3=-6
(3x-2*2x)/(2*3)=-6
(3x-4x)/6=-6
-x/6=-6
-x=-6*6
-x=-36
x=36 ans.
Proof:
36/2+6=2/3*36
18+6=72/3
24=24